Math 042: Exam II Date: April 02 , 2010 Name: Section:

1. Which of the following is the graph of −3x + y = −6?

2. What is the slope of the line given by the equation x = 5?

(a) 1 (b) 0 (c) −5 (d) undefined

2  Exam II  April 02 , 2010

3. Find the slope and the equation of the line that contains the point (−1, 3) and is perpendicular to the line 2x − y = 4. Slope: . Equation: .

4. Find the equation of the line passing through (−2, 5) and (3, −3). Equation: .

5. Solve the following system of equations.

2x − y = 1
−3x + 2y = 5

Your answer: .

6. Graph 3x + 2y < 6.

7. Which of the following is an expansion of (3x − 4y)^{2}?

(a) 9x^{2} + 16y^{2} (b) 9x^{2} − 16y^{2} (c) 9x^{2} − 24xy + 16y^{2} (d) 9x^{2} − 12xy + 16y^{2}

8. Expand (x + 2)^{3}. Your answer: .

4  Exam II  April 02 , 2010

9. x − 49x^{3} = .

10. Factor 7x − 7y − bx + by. Your answer: .

11. Simplify \frac{(2xy^{−3})^{3}}{2x^{2}y^{−2}} . Your answer: .

12. Factor 3x^{2} + 4x − 15. Your answer: .

13. Solve x^{2} − 2x − 8 = 0. Your answer: .

14. Which one of the following is a factor of 8x^{3} − 40x?

(a) 8x^{2} (b) x^{2} − 5 (c) x − 5 (d) x^{2} − 5x

15. (\frac{2}{3})^{−3} = .

(a) −\frac{8}{27} (b) \frac{8}{27} (c) −\frac{27}{8} (d) \frac{27}{8}